find the perimeter of a rectangle 10.53cm wide and 12.04cm long

what is the sum of the measures of the angles in a parallelogram


i would appreciate any help you can give me

Answer by sarah_adam(201)

perimeter of a rectangle : 2(L+W)
where L = length and W = width
Therfore perimeter = 2(10.53+12.04)= 2(22.57) = 45.14 cm
Parallelogram:
A four-sided polygon with two pairs of parallel sides. The sum of the angles of a parallelogram is 360 degrees.